# Vacuum Rabi Splitting and Quantum Fisher Information of a Non-Hermitian Qubit in a Single-Mode Cavity

### v1

@@ II.2 Adiabatic approximation method
A previous study of the PTQRM [17] primarily discussed the adiabatic approximation (AA). In this subsection, we present a simpler alternative to the AA method. Our goal is to establish a practical framework for the further correction and systematic improvement of the AA results.
@@
Within the framework of Eqs. (13) and (14), analytical approximations can be systematically performed. First, as a zeroth-order approximation, we omit the off-diagonal terms and retain only the same $m$-th manifold,
@@
$-D_{m,m}u_{m}+\left(m-g^{2}-i\frac{\epsilon}{2}-E\right)v_{m}$ $=$ $0$ (17)
The nonzero coefficients $u_{m}$ and $v_{m}$ yield the following equation.
@@
where $m=0,1,2...$, the corresponding $m$-th pair of eigenstates is
@@
$\frac{1}{2\eta}i\left(\sqrt{\epsilon^{2}-4\left(D_{m,m}\right)^{2}}+\epsilon\right)u_{m}$ $=$ $v_{m}.$
Set $c_{m}=1,$ we have
@@
It can be confirmed that the wavefunction (21) lacks $\mathcal{PT}$-symmetry, even though the Hamiltonian still commutes with the $\mathcal{P}\mathcal{T}$ operator. This behavior contrasts with the case where the system is in the $\mathcal{PT}$-unbroken phase. As an approximate solution, it can be verified that $\left|m\right\rangle_{\pm}$ must coincide at the EPs where $\epsilon=2\eta$, as given in Eq. (19).
Note that the zeroth-order approximation here is identical to the AA discussed by Lu et al. [17]. The transitions between different manifolds $\left|m\right\rangle_{{A{\pm}}}$ are omitted in this approximation of the wavefunction (15), which is commonly referred to as the adiabatic approximation.
The AA results for the eigenvalues, shown in Fig. 3 for $\epsilon=0.1$ and $\epsilon=0.5$, basically align with the ED results but still deviate in both the $\mathcal{PT}$-unbroken and $\mathcal{P}\mathcal{T}$-broken region, except in the deep-strong coupling regime. There is still room for improving the AA. The key assumption of the AA is that only transitions between the two same manifolds are considered, while transitions between two different manifolds are neglected [45]. To obtain more accurate results, additional transitions must be included. Corrections to the AA method are discussed in the next subsection.
II.3 Corrections to the AA
@@
Figure 3: (colored online) The real part (left panel) of the first four pairs and the imaginary part (right panel) of only the fourth pair of eigenvalues (for clarity) as a function of coupling strength $g$ at resonance $\Delta=1$ for $\epsilon=0.1$ (upper panel) and $\epsilon=0.5$ (lower panel). The black lines represent the ED results, the blue dots show the AA results, and the red dots indicate the CAA results. The eigenvalues and corresponding eigenstates along the dashed vertical line will be used to calculate the dynamics shown in Fig. 4, 5 and 6.
@@
where the first $\mp$ in the second line corresponds to odd and even $m$. The $m$-th pair of eigenvalues can be obtained by selecting the determinant in a $6\times 6$ block form with $n=m,m\pm 1$ in Eqs. (13) and (14), as presented in the Appendix. This determinant leads to a polynomial equation of degree six. An explicit analytical solution cannot be obtained due to the Abel-Ruffini theorem, which asserts that no general solution in radicals exists for polynomial equations of degree five or higher with arbitrary coefficients [46]. There are six groups of roots, but only two are required. The CAA results should match the AA results from Eq. (18). This allows us to unambiguously identify the $m$-th pair of eigenvalues in the CAA from the full set of solutions. A simple selection rule is outlined in the Appendix.
@@
Further corrections, such as the inclusion of five sets of coefficients, $\left(u_{m\pm i},v_{m\pm i}\right),i=0,1,2$, can also be implemented in a similar manner. In principle, these corrections can be applied systematically, step by step, ultimately converging to the exact solutions outlined in Subsection A. However, if a qualitative understanding of the physics is achieved within the CAA, additional corrections may be unnecessary in the analytical sense, as shown in the next section, where the CAA will also be used to calculate the qubit dynamics and study the VRS of the PTQRM.
@@ III Dynamics and vacuum Rabi splitting
One of the notable features in atom-cavity coupling system is the VRS, as described in [47, 48, 49]. The atom is typically pumped from its ground state to the excited state, and then decays to the ground state of the whole system. The resulting emission spectrum displays two peaks of equal height at resonance, characteristic of VRS in the Jaynes-Cummings model [26].
In this section, we calculate the time evolution of the qubit population difference, $\left\langle\sigma_{z}\right\rangle$, in the PTQRM using three methods: the CAA, AA, and exact solution, either ED or through the exact eigensolution obtained via the BOA. We use $\left|\psi_{n}^{R}\right\rangle$ and $\left|\psi_{n}^{L}\right\rangle$ to denote the right and left eigenstates, the eigenstates of $H$ and $H^{\dagger}$, respectively. The normalization conditions are $\left\langle\psi_{n}^{R}\mid\psi_{n}^{R}\right\rangle=1$ and $\left\langle\psi_{m}^{L}\mid\psi_{n}^{R}\right\rangle=$ $\delta_{m,n}$, while the right and left eigenstates with different eigenvalues are non-orthogonal. That is $\left\langle\psi_{m}^{R(L)}\mid\psi_{n}^{R(L)}\right\rangle\neq 0$ for $m\neq n$, due to the non-Hermitian nature  [50]. The time-dependent wave function of the PTQRM, starting from $|\Psi(0)\rangle$, can be expressed as
@@
The initial state is chosen as the atomic upper level and the photon vacuum state in the PTQRM, similar to the conventional study of the VRS. The dynamical evolution at resonance ($\Delta=1$) is calculated using the CAA for two imaginary biases, $\epsilon=0.1$ and $\epsilon=0.5$, and three typical coupling strengths: $g=0.1$, $0.75$, and $1.25$. These coupling strengths range from the ultrastrong to deep-strong coupling regimes. The results (red line) for the dynamics are respectively shown in the left panel of Figs.4,  5, and  6, while the emission spectrum of the atomic population evolution is shown in the right panel of these three figures. Results from the ED (black dashed line) and AA (blue line) are also presented for comparison.
@@
Ultrastrong coupling regime ($g=0.2$):- As shown in Fig. 4, the CAA results perfectly agree with the exact ones in both atomic dynamics and the emission spectrum. Oscillations with two main frequency are observed in the PTQRM, similar to the VRS in its Hermitian counterpart. This demonstrates that $\mathcal{PT}$-symmetry is preserved in the local phase during the early stages of the time evolution. In contrast, a oscillation with a single frequency is exhibited in the AA. Note that subharmonic peaks in the high frequency region of the emission spectrum result from the imperfect sinusoidal oscillations and do not indicate oscillations at new frequencies.
Figure 4: (colored online) Time evolution of $<\sigma_{z}>$ (left panel) and the corresponding emission spectrum (right panel) at resonance ($\Delta=1$) for the ultrastrong coupling ($g=0.2$) with $\epsilon=0.1$ (upper panel) and $\epsilon=0.5$ (lower panel). The peak positions are also shown in the left panel.
Interestingly, the positions of the two peaks in the CAA coincide with the crossing points of the vertical dashed line and the CAA eigenvalues in the left panel of Fig. 3, corresponding to the energies for the first and second excited states relative to the ground state energy, $E^{CAA}_{1}-E^{CAA}_{GS}$ and $E^{CAA}_{2}-E^{CAA}_{GS}$, respectively. In contrast, the position of the single peak in the AA reflects the difference between the first excited state energy ($E_{+}$) and the ground state energy ($E_{-}$), as given in Eq.(18), and is also shown in the left panel of Fig. 3. This occurs because, in the CAA, the nearest-neighboring manifold is considered, allowing transitions from both the first and second excited states to the ground state. In contrast, only the transition between the same manifold is taken into account in the AA, so the second excited state is excluded.
Figure 5: (colored online) Time evolution of $<\sigma_{z}>$ (left panel) and the corresponding emission spectrum (right panel) at resonance ($\Delta=1$) for near deep-strong coupling ($g=0.75$) with $\epsilon=0.1$ (upper panel) and $\epsilon=0.5$ (lower panel).
Near deep-strong coupling regime ($g=0.75$):- As shown in the upper panel of Fig. 5 for both biases, the population from both the CAA and exact solutions oscillates with more than two frequencies, indicating multiple VRS. This is similar to the multiple Rabi oscillations observed in the ultrastrong coupling regime of its Hermitian counterpart without biases [51] and in systems with ultrastrong molecular vibrational coupling [52]. However, the frequency spectrum in the AA shows only one peak for $\epsilon=0.1$, and the $<\sigma_{z}(t)>$ curve tracks only the envelope of the multiple oscillations observed in the exact solutions, erasing the details of the fast oscillations. No peak at finite frequency is observed in the AA for large bias $\epsilon=0.5$.
@@
Figure 6: (colored online) Time evolution of $<\sigma_{z}>$ (left panel) and the corresponding emission spectrum (right panel) at resonance ($\Delta=1$) for the deep-strong coupling ($g=1.25$) with $\epsilon=0.1$ (upper panel) and $\epsilon=0.5$ (lower panel).
Deep-strong coupling regime $g=1.25$:- In this coupling regime, the atomic population difference never becomes negative, indicating that the atom predominantly stays in the upper level. The atomic population oscillates also with multiple frequencies. The CAA matches the exact solution, while the AA only captures the general envelope of the true dynamics, eliminating all the fast oscillations.
@@
In this non-Hermitian $\mathcal{PT}$-symmetric system, the atom tends to return to its original state, accompanied by an infinite number of photons [17]. This feature is clearly visible in Figs. 4 (c),  5, and  6 during the final time window. For the small bias and ultrastrong coupling regime, we calculated the long-time evolution, continuing from the data in Figs. 4 (a), and observed similar behavior after a prolonged oscillation (not shown here) . This represents the final outcome, likely due to the balance of gain and loss in the non-Hermitian system with $\mathcal{PT}$ symmetry.
@@
In quantum parameter estimation theory, the QFI describes the minimum error in extracting an unknown parameter from a given quantum state. The optimal measurement accuracy is governed by the quantum Cramér-Rao bound  [55, 56, 29]. It can be proven that the expressions for both the traditional QFI and the parameter estimates of the quantum Cramér-Rao bound still hold for non-Hermitian Hamiltonians under stationary-states [29, 57]. Thus, for the PTQRM, the QFI for the pure states is defined as
@@
where $\lambda$ represents the experimental parameter to be estimated. In the present study, we select $\lambda=g$, $|\psi(g)\rangle$ is one eigenstate and $|\psi^{\prime}(g)\rangle$ is its first-order derivative with respect to the coupling strength $g$. As mentioned previously, for a non-Hermitian Hamiltonian ($H^{{\dagger}}\neq H$), the eigenstates are typically non-orthogonal. Under the stationary state condition, the normalization condition of the eigenstates holds, and it can be proven that the QFI still satisfies Eq. (28). We have verified that $|\psi(g)\rangle$ can be selected either the right or the left eigenstates of the PTQRM, the QFI remains unchanged.
Figure 8 (a) illustrates the QFIs in the ground state for both the Hermitian QRM and PTQRM, calculated using the wavefunction derived from the G-function technique or the ED. The results are shown for $\Delta=0.5$ and various bias parameters ($\epsilon=0,0.2,0.4$). The QFI in the PTQRM are noticeably higher than those in the Hermitian QRMs, both in the symmetric case ($\epsilon=0$) and the asymmetric case ($\epsilon\neq 0$). Near the EPs $g=0.6828$ for $\epsilon=0.2$ and $g=0.3358$ for $\epsilon=0.4$, the QFI in the PTQRM exhibits a sharp peak. This indicates that the PT-symmetric non-Hermitian system is highly sensitive to perturbations in the estimated parameter near the EP. This effect is known as $\mathcal{PT}$-enhanced quantum sensing [32, 34, 35]. In the Hermitian QRM, the QFI shows a shallow maximum. The Hermitian QRM shows only a crossover, without a phase transition exhibiting a singularity. In contrast, the EPs of the PTQRM mark the transition from the $\mathcal{PT}$-unbroken phase to the $\mathcal{P}\mathcal{T}$-broken phase. Figure 8 (b) shows that the QFI in the fourth excited state near the first EP is significantly higher than that in the ground state for the same model parameters. This implies that using the higher excited state can improve the estimation accuracy of unknown parameters.
Since the PTQRM can be viewed as a NH-TLS coupled to a quantized cavity. A natural question arises: can quantum sensing for the NH-TLS be enhanced through coupling with the cavity. To address this issue, we calculate the QFI for both the NH-TLS and PTQRM, where the former is a component of the latter. To provide a clear comparison, we derive the analytical QFI for the NH-TLS and PTQRM in the AA below,
$F_{\epsilon}^{NH-TLS}=\left\{\begin{array}[]{lll}\frac{1}{4\left(\Delta^{2}-\epsilon^{2}\right)},&\epsilon\leqslant\Delta\\ \frac{\Delta^{2}}{4\epsilon^{2}\left(\epsilon^{2}-\Delta^{2}\right)},&\epsilon>\Delta\end{array}\right.$ (29)
@@
where EPs are given by $\epsilon_{EP}=\Delta$ for the NH-TLS and $\Delta e^{-2g^{2}}$ for the PTQRM.
In the $\mathcal{PT}$-unbroken regime ($\epsilon\leqslant\epsilon_{EP}$), let $\epsilon$ deviate slightly from the EP by a small amount $\delta$; thus, the QFI becomes
@@
$F_{\epsilon=\Delta-\delta}^{NH-TLS}=\left(8\Delta\delta\right)^{-1},$
@@
for the NH-TLS. Here, higher-order terms in small $\delta$ is neglected. The QFI is finite as long as $\delta\neq 0$. It is evident that the QFI in the PTQRM is higher than that in the NH-TLS, with an enhanced factor of $e^{2g^{2}}$. At weak coupling, the enhanced factor can be neglected; however, it becomes more pronounced as the coupling strength increases. Clearly, these findings also apply to the $\mathcal{PT}$-broken regime.
@@
The numerically exact QFI (solid colored line) and the analytical QFI in the AA (dashed line) for the PTQRM are shown in Fig. 8 (c), along with the QFI for the NH-TLS (black solid line) for comparison. At weak coupling ($g=0.1$), the QFI remains nearly unaffected by the cavity coupling but increases as the coupling strength grows. In the deep-strong coupling regime ($g=1.0$), the QFI is significantly higher than that in the NH-TLS. This can be more clear if we normalize the bias value to the peak position (also at EP), as shown in Fig. 8 (d). Strong qubit-cavity coupling offers a novel method to enhance quantum sensing in the extensively studied NH-TLS. Note that the QFI calculated by the AA is quite accurate because only the same eigenstates are used.

### v2

@@ II.2 Adiabatic approximation method
A previous study of the PTQRM [17] primarily discussed the AA. In this section, we present a simpler alternative to the AA method. Our goal is to establish a practical framework for the further correction and systematic improvement of the AA results.
@@
For the truly exact solution, the summation in the second term of Eqs. (14) and (15) must be infinite. Physically, tunneling between all the $m$th manifolds must be considered. The overlap $D_{mm}$ within the same $m$th manifold is larger than the overlaps $D_{mn}$ between different manifolds ($n\neq m$). Based on these observations, a simple preliminary analytical approach is to retain only the transitions between states within the same manifold.This is actually the adiabatic approximation introduced by Irish et al. for the Hermitian QRM [47] and later extended to the two-qubit QRM [48]. Further corrections can be made by including the transitions between different manifolds. Therefore, within the framework of Eqs. (14) and (15), analytical approximations can be systematically performed. First, as a zero-order approximation, we omit the off-diagonal terms and retain only the terms within the same $m$th manifold,
@@
$-D_{m,m}u_{m}+\left(m-g^{2}-i\frac{\epsilon}{2}-E\right)v_{m}$ $=$ $0.$ (18)
The nonzero coefficients $u_{m}$ and $v_{m}$ yield the following equation:
@@
where $m=0,1,2$..., and the corresponding $m$-th pair of eigenstates is
@@
$\frac{1}{2{D_{m,m}}}i\left(\sqrt{\epsilon^{2}-4\left(D_{m,m}\right)^{2}}+\epsilon\right)u_{m}$ $=$ $v_{m}.$
Setting $c_{m}=1$, we have
@@
It can be confirmed that wavefunction (22) lacks $\mathcal{PT}$symmetry, even though the Hamiltonian still commutes with the $\mathcal{P}\mathcal{T}$ operator. This behavior contrasts with the case where the system is in the $\mathcal{PT}$-unbroken phase. As an approximate solution, it can be verified that $\left|m\right\rangle_{\pm}$ must coincide at the EPs where $\epsilon=2D_{m,m}$, as given in Eq. (19).
Note that the zero-order approximation here is identical to the AA discussed by Lu et al. [17]. The transitions between different manifolds $\left|m\right\rangle_{{A{\pm}}}$ are omitted in this approximation of the wavefunction (16), which is commonly referred to as the adiabatic approximation.
The AA results for the eigenvalues, shown in Fig. 3 for $\epsilon=0.1$ and $\epsilon=0.5$, basically align with the ED results but still deviate in both the $\mathcal{PT}$-unbroken and $\mathcal{P}\mathcal{T}$-broken region, except in the deep strong-coupling regime. There is still room for improving the AA. The key assumption of the AA is that only transitions within the same manifolds are considered, while transitions between different manifolds are neglected [49], rendering the AA approach trivial. To obtain more accurate results, additional transitions must be included. Corrections to the AA method are discussed in the next section.
II.3 Corrections to the adiabatic approximation
@@
Figure 3: The real part (left) of the first four pairs and the imaginary part (right) of only the fourth pair of eigenvalues (for clarity) as a function of coupling strength $g$ at resonance $\Delta=1$ for $\epsilon=0.1$ (upper panel) and $\epsilon=0.5$ (lower panel). The black lines represent the ED results, the blue dots show the AA results, and the red dots indicate the CAA results. The eigenvalues and corresponding eigenstates along the dashed vertical line will be used to calculate the dynamics shown in Fig. 4, 5 and 6.
@@
where the first $\mp$ in the second line corresponds to odd and even $m$. The $m$-th pair of eigenvalues can be obtained by selecting the determinant in a $6\times 6$ block form with $n=m,m\pm 1$ in Eqs. (14) and (15), as presented in the Appendix. This determinant leads to a polynomial equation of degree $6$. An explicit analytical solution cannot be obtained due to the Abel-Ruffini theorem, which asserts that no general solution in radicals exists for polynomial equations of degree $5$ or higher with arbitrary coefficients [50]. There are six groups of roots, but only two are required. The CAA results should match the AA results from Eq. (19). This allows us to unambiguously identify the $m$-th pair of eigenvalues in the CAA from the full set of solutions. A simple selection rule is outlined in the Appendix.
@@
Further corrections, such as the inclusion of five sets of coefficients, $\left(u_{m\pm i},v_{m\pm i}\right),i=0,1,2$, can also be implemented in a similar manner. In principle, these corrections can be applied systematically, step by step, ultimately converging to the exact solutions outlined in Sec. II A. However, if a qualitative understanding of the physics is achieved within the CAA, additional corrections may be unnecessary in the analytical sense, as shown in the next section, where the CAA will also be used to calculate the qubit dynamics and study the VRS of the PTQRM.
@@ III Dynamics and vacuum Rabi splitting
One of the notable features in atom-cavity coupling system is the VRS, as described in Refs.[51, 52, 53]. The atom is typically pumped from its ground state to the excited state, and then decays to the ground state of the whole system. The resulting emission spectrum displays two peaks of equal height at resonance, characteristic of VRS in the Jaynes-Cummings model [26].
In this section, we calculate the time evolution of the qubit population difference, $\left\langle\sigma_{z}\right\rangle$, in the PTQRM using three methods: the CAA, the AA, and the exact solution, either ED or through the exact eigensolution obtained via the BOA. We use $\left|\psi_{n}^{R}\right\rangle$ and $\left|\psi_{n}^{L}\right\rangle$ to denote the right and left eigenstates, the eigenstates of $H$ and $H^{\dagger}$, respectively. The normalization conditions are $\left\langle\psi_{n}^{R}\mid\psi_{n}^{R}\right\rangle=1$ and $\left\langle\psi_{m}^{L}\mid\psi_{n}^{R}\right\rangle=$ $\delta_{m,n}$, while the right and left eigenstates with different eigenvalues are nonorthogonal. That is $\left\langle\psi_{m}^{R(L)}\mid\psi_{n}^{R(L)}\right\rangle\neq 0$ for $m\neq n$, due to the non-Hermitian nature  [54]. The time-dependent wave function of the PTQRM, starting from $|\Psi(0)\rangle$, can be expressed as
@@
The initial state is chosen as the atomic upper level and the photon vacuum state in the PTQRM, similar to the conventional study of the VRS. The dynamical evolution at resonance ($\Delta=1$) is calculated using the CAA for two imaginary biases, $\epsilon=0.1$ and $\epsilon=0.5$, and three typical coupling strengths: $g=0.1$, $0.75$, and $1.25$. These coupling strengths range from the ultrastrong to deep-strong coupling regimes. The results (red line) for the dynamics are respectively shown in the left-panel of Figs.4,  5, and  6, while the emission spectrum of the atomic population evolution is shown in the right-panel of these three figures. Results from the ED ( dashed black line) and AA (blue line) are also presented for comparison.
@@
Ultrastrong coupling regime ($g=0.2$). As shown in Fig. 4, the CAA results perfectly agree with the exact ones in both atomic dynamics and the emission spectrum. Oscillations with two main frequencies are observed in the PTQRM, similar to the VRS in its Hermitian counterpart. This demonstrates that $\mathcal{PT}$-symmetry is preserved in the local phase during the early stages of the time evolution. In contrast, a oscillation with a single frequency is exhibited in the AA. Note that subharmonic peaks in the high frequency region of the emission spectrum result from the imperfect sinusoidal oscillations and do not indicate oscillations at new frequencies.
Figure 4: Time evolution of $\sigma_{z}>$ (left) and the corresponding emission spectrum (right) at resonance ($\Delta=1$) for the ultrastrong coupling ($g=0.2$) with $\epsilon=0.1$ (upper panel) and $\epsilon=0.5$ (lower panel). The peak positions are also shown in the left-hand panel.
Interestingly, the positions of the two peaks in the CAA coincide with the crossing points of the vertical dashed line and the CAA eigenvalues in the left-hand panel of Fig. 3, corresponding to the energies for the first and second excited states relative to the ground state energy, $E^{CAA}_{1}-E^{CAA}_{GS}$ and $E^{CAA}_{2}-E^{CAA}_{GS}$, respectively. In contrast, the position of the single peak in the AA reflects the difference between the first excited state energy ($E_{+}$) and the ground state energy ($E_{-}$), as given in Eq.(19), and is also shown in the left-hand panel of Fig. 3. This occurs because, in the CAA, the nearest-neighboring manifold is considered, allowing transitions from both the first and second excited states to the ground state. In contrast, only the transition between the same manifold is taken into account in the AA, so the second excited state is excluded.
Figure 5: Time evolution of $\sigma_{z}>$ (left) and the corresponding emission spectrum (right) at resonance ($\Delta=1$) for near deep strong coupling ($g=0.75$) with $\epsilon=0.1$ (upper panel) and $\epsilon=0.5$ (lower panel).
Near deep-strong coupling regime ($g=0.75$).- As shown in the upper panel of Fig. 5 for both biases, the population from both the CAA and exact solutions oscillates with more than two frequencies, indicating multiple VRS. This is similar to the multiple Rabi oscillations observed in the ultrastrong coupling regime of its Hermitian counterpart without biases [55] and in systems with ultrastrong molecular vibrational coupling [56]. However, the frequency spectrum in the AA shows only one peak for $\epsilon=0.1$, and the $\sigma_{z}(t)>$ curve tracks only the envelope of the multiple oscillations observed in the exact solutions, erasing the details of the fast oscillations. No peak at finite frequency is observed in the AA for large bias $\epsilon=0.5$.
@@
Figure 6: Time evolution of $\sigma_{z}>$ (left) and the corresponding emission spectrum (right) at resonance ($\Delta=1$) for the deep strong coupling ($g=1.25$) with $\epsilon=0.1$ (upper panel) and $\epsilon=0.5$ (lower panel).
Deep-strong coupling regime $g=1.25$.- In this coupling regime, the atomic population difference never becomes negative, indicating that the atom predominantly stays in the upper level. The atomic population oscillates also with multiple frequencies. The CAA matches the exact solution, while the AA only captures the general envelope of the true dynamics, eliminating all the fast oscillations.
@@
In this non-Hermitian $\mathcal{PT}$-symmetric system, the atom tends to return to its original state, accompanied by an infinite number of photons [17]. This feature is clearly visible in Figs. 4(c),  5, and  6 during the final time window. For the small bias and ultrastrong coupling regime, we calculated the long-time evolution, continuing from the data in Fig. 4(a), and observed similar behavior after a prolonged oscillation (not shown here) . This represents the final outcome, likely due to the balance of gain and loss in the non-Hermitian system with $\mathcal{PT}$ symmetry.
@@
In quantum parameter estimation theory, the QFI describes the minimum error in extracting an unknown parameter from a given quantum state. The optimal measurement accuracy is governed by the quantum Cramér-Rao bound  [58, 59, 29]. It can be proven that the expressions for both the traditional QFI and the parameter estimates of the quantum Cramér-Rao bound still hold for non-Hermitian Hamiltonians under stationary-states [29, 60, 61]. Thus, for the PTQRM, the QFI for the pure states is defined as
@@
where $\lambda$ represents the experimental parameter to be estimated. In the present study, we select $\lambda=g$, $|\psi(g)\rangle$ is one eigenstate and $|\psi^{\prime}(g)\rangle$ is its first-order derivative with respect to the coupling strength $g$. As mentioned previously, for a non-Hermitian Hamiltonian ($H^{{\dagger}}\neq H$), the eigenstates are typically nonorthogonal. Under the stationary state condition, the normalization condition of the eigenstates holds, and it can be proven that the QFI still satisfies Eq. (29). We have verified that $|\psi(g)\rangle$ can be selected either the right or the left eigenstates of the PTQRM, the QFI remains unchanged.
Both the static  [62, 63] and dynamic  [64, 65] approaches are commonly used in the critical quantum metrology protocol. In this work, we focus on measuring the QFI in equilibrium, a method that falls under the static approach. Figure 8 (a) illustrates the difference in QFIs between the PTQRM and the Hermitian QRM with $\Delta=0.5$ in the ground state as a function of coupling strength $g$ and $\epsilon$, calculated using either the wavefunction derived from the $G$ function technique or the ED. Throughout the whole parameter regime plotted (up to the deep strong coupling $g=1.5$), the QFI of the PTQRM exhibits a significant enhancement near the EPs compared to the Hermitian QRM. Figure 8 (c) shows the results for various bias parameters ($\epsilon=0,0.2,0.4$). The QFI in the PTQRM is significantly higher than those in the Hermitian QRMs, both in the symmetric case ($\epsilon=0$) and the asymmetric case ($\epsilon\neq 0$). Near the EPs $g=0.6828$ for $\epsilon=0.2$ and $g=0.3358$ for $\epsilon=0.4$, the QFI in the PTQRM exhibits a sharp peak. This suggests that the $\mathcal{PT}$-symmetric non-Hermitian system is highly sensitive to perturbations in the estimated parameter near the EP. This effect is known as $\mathcal{PT}$-enhanced quantum sensing  [32, 34, 35]. In the Hermitian QRM, the QFI shows a shallow maximum. The Hermitian QRM shows only a crossover, without a phase transition exhibiting a singularity. In contrast, the EPs of the PTQRM mark the transition from the $\mathcal{PT}$-unbroken phase to the $\mathcal{P}\mathcal{T}$-broken phase. Figure 8 (d) shows that the QFI in the fourth excited state near the first EP is significantly higher than that in the ground state for the same model parameters. This implies that using the higher excited state can improve the estimation accuracy of unknown parameters.
Since the PTQRM can be viewed as a NHTLS coupled to a quantized cavity, a natural question arises: Can quantum sensing for the NHTLS be enhanced through coupling with the cavity. To address this issue, we calculate the QFI for both the NHTLS and PTQRM, where the former is a component of the latter. To provide a clear comparison, we derive the analytical QFI for the NHTLS and PTQRM in the AA below,
$F_{\epsilon}^{NHTLS}=\left\{\begin{array}[]{lll}\frac{1}{4\left(\Delta^{2}-\epsilon^{2}\right)},&\epsilon\leqslant\Delta\\ \frac{\Delta^{2}}{4\epsilon^{2}\left(\epsilon^{2}-\Delta^{2}\right)},&\epsilon>\Delta\end{array}\right$. (30)
@@
where EPs are given by $\epsilon_{EP}=\Delta$ for the NHTLS and $\Delta e^{-2g^{2}}$ for the PTQRM.
In the $\mathcal{PT}$-unbroken regime ($\epsilon\leqslant\epsilon_{EP}$), let bias $\epsilon$ deviate slightly from the EP by a small amount $\delta$; thus, the QFI becomes
@@
$F_{\epsilon=\Delta-\delta}^{NHTLS}=\left(8\Delta\delta\right)^{-1},$
@@
for the NHTLS. Here, higher-order terms in small $\delta$ are neglected. The QFI is finite as long as $\delta\neq 0$. It is evident that the QFI in the PTQRM is higher than that in the NHTLS, with an enhanced factor of $e^{2g^{2}}$. At weak coupling, the enhanced factor can be neglected; however, it becomes more pronounced as the coupling strength increases. Clearly, these findings also apply to the $\mathcal{PT}$-broken regime.
@@
Figure 8(b) shows the difference in QFIs between the PTQRM and the NHTLS in the ground state as a function of coupling strength $g$ and bias $\epsilon$. Both the PTQRM and NHTLS show a significant enhancement near the EPs. The numerically exact QFI (solid colored line) and the analytical QFI in the AA (dashed line) for the PTQRM are shown in Fig. 8(d), along with the QFI for the NHTLS ( solid black line) for comparison. The bias value is normalized with respect to the corresponding EPs. At weak coupling ($g=0.1$), the QFI is nearly unaffected by the cavity coupling but increases as the coupling strength increases. In the deep strong-coupling regime ($g=1.0$), the QFI is substantially higher than that in the NHTLS. Strong qubit-cavity coupling offers an alternative approach to enhance quantum sensing in the extensively studied NHTLS. Note that the QFI calculated by the AA is quite accurate because only the same eigenstates are used.